# Vipera: Blending Visual and LLM-Driven Guidance for Systematic Auditing of Text-to-Image Generative AI

### v1

@@ Abstract.
Despite their increasing capabilities, text-to-image generative AI systems are known to produce biased, offensive, and otherwise problematic outputs. While recent advancements have supported testing and auditing of generative AI, existing auditing methods still face challenges in supporting effectively explore the vast space of AI-generated outputs in a structured way. To address this gap, we conducted formative studies with five AI auditors and synthesized five design goals for supporting systematic AI audits. Based on these insights, we developed Vipera, an interactive auditing interface that employs multiple visual cues including a scene graph to facilitate image sensemaking and inspire auditors to explore and hierarchically organize the auditing criteria. Additionally, Vipera leverages LLM-powered suggestions to facilitate exploration of unexplored auditing directions. Through a controlled experiment with 24 participants experienced in AI auditing, we demonstrate Vipera’s effectiveness in helping auditors navigate large AI output spaces and organize their analyses while engaging with diverse criteria. [^1] The work is in submission.
AI Auditing, Generative Text-To-Image Models, Visual Analytics
$c$s: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interactive systems and tools
@@ 1. Introduction
Generative text-to-image (T2I) models are gaining popularity for their ability to enhance the efficiency and expressiveness of creative activities. However, the associated risks are significant; these models can produce images that may perpetuate biases, cause offense, or disseminate misleading information (Naik and Nushi, 2023; Qi et al., 2023). To surface and address these issues, researchers and practitioners have turned to AI auditing—a process of repeatedly testing an algorithm with inputs and observing the corresponding outputs—in order to better understand its behavior and potential external impacts. (Sandvig et al., 2014; Metaxa et al., 2021; Birhane et al., 2024).
@@
Recent research in HCI and responsible AI (RAI) has explored tools and processes to support diverse auditors in AI audits (Lam et al., 2022; Shen et al., 2021; Cabrera et al., 2021a; Deng et al., 2023a; Maldaner et al., 2024; Claire et al., 2024). However, auditing generative T2I models in a systematic and scaled way remains challenging. To start with, manual evaluation of these systems can be time-consuming, often limiting assessments to a small number of prompts or outputs, which can hardly uncover the comprehensive issues of the system. Developers often face challenges in exploring potentially productive auditing directions and gathering scaled data for each identified issue (Lam et al., 2022; Deng et al., 2025a). Reviewing and aggregating audit reports also remains a demanding task (Cabrera et al., 2021a; Deng et al., 2023a; Ojewale et al., 2024)
@@
To tackle these challenges, automated approaches have also been proposed, often involving AI-supported output labeling and applied to textual models (Shankar et al., 2024; Rastogi et al., 2023a; Arawjo et al., 2024). However, two primary obstacles persist when it comes to auditing T2I models. First, images encompass a broad range of semantics, resulting in a vast auditing space that is difficult to characterize with a concise list of criteria as often used for evaluating texts (Suh et al., 2024). For example, an image may feature multiple characters, each of whom can be evaluated based on the clothing, and the clothing can be further assessed by its cultural nature and relationship to the character’s profession. To systematically explore this space, it is crucial to keep auditors aware of the unknown unknowns - criteria that are yet to be explored (Kiela et al., 2021; Deng et al., 2025a). Additionally, our observations from a formative study with five auditors indicate that many auditors often rely on intuition or personal experience during the auditing process without a formal approach. There is a lack of structured methodologies for navigating and exploring this extensive auditing space.
@@
In this paper, we propose Vipera [^2] Vipera stands for ”Visual Intelligence-Promoted End User Auditing”, an interactive system for streamlining and enhancing the systematicness of large-scale T2I model auditing. Vipera facilitates structured multi-faceted analysis through the visual cue of scene graph (Johnson et al., 2015), where the auditing criteria are organized hierarchically, coordinated with visual statistics, and associated with scenic semantics of images. Additionally, Vipera incorporates LLM-powered auditing suggestions to uncover new avenues for analysis and facilitate exploration within the auditing space. In particular, it provides audit analysis support by highlighting differences between images to suggest new criteria, and prompt suggestions to inspire exploration of new topics, with optional user-customizable keywords for targeted guidance. To support auditors in documenting and reporting their findings, Vipera allows users to bookmark visual evidence immediately upon getting insights, incorporates a dedicated note view for structured note construction, and supports LLM-powered note completion for better efficiency.
@@
Through a controlled user study with 20 general auditors and 4 expert auditors, we demonstrate Vipera’s effectiveness in helping auditors navigate the large auditing space and organize their analyses while engaging with diverse criteria. In particular, we discovered that the blended guidance prompted participants to explore more prompts, images, and auditing criteria compared with conditions without guidance (see Section 6.1.2). We also show that visual and LLM-driven guidance are highly complementary, and blending them can reduce user workload, improve performance, and encourage a more systematic auditing process (see Section 6.2.3). Furthermore, we have revealed several interesting patterns in their auditing process that serve as the basis for future personalized auditing tools (see Section 6.2.4). For example, we observed breadth-oriented and depth-oriented patterns in both the creation of prompts and criteria, suggesting a need for personalized auditing guidance.
@@
  • An empirical study on 24 participants with various background in auditing that evaluates the system’s usability as well as the distinct and combined effects of these guidance modalities, demonstrating that their integration leads to more systematic and thorough auditing with reduced cognitive load compared to three ablated versions of the system.
@@
Figure 1. The Vipera interface. (A) The input box for creating prompts and specifying the number of images. (B) The analysis view for interactive auditing. Left: The prompts and AI-powered suggestions of auditing criteria and prompts. Right: The generated images and an interactive scene graph summarizing the image semantics to inspire users and guide the auditing process. (C) The note view for composing an auditing report. :
@@ 2.1. Auditing generative AI (at scale)
Generative AI (GenAI) systems have sparked increasing societal concerns due to potential harmful behaviors, such as social biases and violence (Diberardino et al., 2024). This has led to increased focus on detecting and mitigating these issues through benchmarks like SafetyBench (Zhang et al., 2023) and AgentHarm (Andriushchenko et al., 2024). However, the limited diversity of inputs and contexts in these benchmarks hampers real-world performance assessments (Kiela et al., 2021).
@@
AI audits have gained prominece as a method for uncovering biased, discriminatory, or otherwise harmful behaviors in algorithmic systems (Noble, 2018; Asplund et al., 2020; Sweeney, 2013; Prates et al., 2020; Buolamwini and Gebru, 2018; Hannak et al., 2014; Sandvig et al., 2014; Metaxa et al., 2021; Birhane et al., 2024; Shen et al., 2022). More recent research have demonstrated the value of conducting AI audits to ensure more safe and responsible GenAI, often even involving general end users in the auditing process, as they can uncover overlooked cases and provide insights (Shen et al., 2021; Lam et al., 2022; Cabrera et al., 2021b; Deng et al., 2025a). For example, Mack et al. engaged 25 people with disabilities to review and reflect on images generated by T2I systems they uncovered several nuanced societal stereotypes that extended beyond existing taxonomies, such as “perpetuating broader narratives in society around disabled people as primarily using wheelchairs, being sad and lonely, incapable, and inactive” (Mack et al., 2024). Similarly, Shelby et al. conducted workshops with 15 cross-cultural artists to synthesize folk theories related to the use of T2I models, their potential harms, and harm reduction strategies perceived by artists (Shelby et al., 2024). Finally, Deng et al. created WeAudit system to engage 45 users in individually and collectively auditing GenAI, with scaffolds to suppport users in providing actionable insights to AI developers (Deng et al., 2025a).
@@
Prior research has primarily engaged people in auditing limited sets of GenAI outputs, while acknowledging the needs for better toolings for scalable, systemtic auditing (Shelby et al., 2024; Deng et al., 2025a). However, research in HCI show that crowdsourced auditing remains challenging and costly due to data scale and often narrow evalaution criteria (Cabrera et al., 2021a; Deng et al., 2023a). To this end, recent studies have begun to focus on scalable auditing of generative AI, primarily addressing two key challenges. First, evaluating large numbers of outputs within limited timeframes: here, human-in-the-loop auditing combines large language models for automatic evaluation with user-facing summaries (Zheng et al., 2024; Shankar et al., 2024; Rastogi et al., 2023b), and annotations and visualizations have been introduced to support interpretation (Gero et al., 2024; Deng et al., 2025b). Second, early auditing tools often relied on rigid quantitative metrics, but recent efforts allow users to define their own criteria (Kim et al., 2024).
@@
However, aforementioned work on improving systematic auditing has mainly focused on auditing text-to-text models. Extending them to text-to-image models introduces new challenges: Images contain rich semantic information (e.g., objects and styles) that can lead to a diverse range of auditing criteria, yet there is a notable lack of tools to assist auditors in interpreting auditing results for iterative assessments in the image context (Pettersson, 1993). To fill this gap, our work introduces Vipera, which integrates visual cues and AI-driven auditing suggestions, to support structured and systematic exploration of auditing criteria in the image domain. In next section, we exand on the unique challenges of auditing image context.
@@ 2.2. Visual analytics for sensemaking image collection
Sensemaking of large image collections plays a key role in auditing T2I models at scale, and visual analytics approaches have proven effective in this context (Afzal et al., 2023). One common method involves clustering images based on pixel or vector embeddings for visualization, which supports various analytical tasks such as search and exploration, comparison, and visual summarization (Zahálka et al., 2020; Schmidt et al., 2013; Pan et al., 2019). To incorporate image semantics alongside visual features, recent works identify semantic objects within images and extract concise textual representations—such as keywords, descriptions, or captions. These text analysis results facilitate tasks like semantic categorization and pattern mining (Xie et al., 2018; Yee et al., 2003; Li et al., 2024). However, characterizing the relationships between objects using plain text can be challenging. To address this, scene graphs have emerged as a more expressive representation, where nodes denote semantic objects and edges illustrate their relationships. This representation enhances detailed understanding of image collections and improves analysis performance (Fang et al., 2017; Johnson et al., 2015).
With the advent of generative T2I models, a recently emerged task is to visualize the relationship of between prompts and corresponding images. For instance, DreamSheets organizes prompts and images in a spreadsheet layout to facilitate comparison (Almeda et al., 2024). Similarly, PrompTHis introduces an image variant graph to illustrate the semantic transitions of image clusters as users create new prompts (Guo et al., 2024). However, these approaches often struggle to support sensemaking in large-scale image sets, where variations among outputs for a single prompt can lead to overly condensed visual clusters.
Drawing from prior work and situating it in the context of AI auditing, Vipera first embeds bar charts within a tree-based scene graph, revealing distributional patterns for images related to individual prompts and allowing for comparisons across prompts. Besides, Vipera leverages the scene graph as a visual aid to foster systematic and structured AI auditing while providing inspiration to auditors. Additionally, Vipera incorporates various view coordinations to facilitate navigation between auditing results and raw images for enhanced sensemaking. Our work further contributes by evaluating these visualizations in the context of systematic AI auditing, offering critical insights for developing more effective auditing tools for text-to-image systems.
@@ 3.1.1. Study setup
Prototypes. Following parallel prototyping (Dow et al., 2010), we designed two prototypes for the formative study, referred to as ViperaBase and Baseline in the following discussions. As illustrated in Figure 2, ViperaBase features a scene graph view, i.e., the zoomable node-link diagram, alongside a prompt input box and image view, whereas the Baseline includes all views except the scene graph view. In this view, blue nodes represent objects, and green nodes indicate their attributes, with labeled edges illustrating relationships. The size of each node reflects the number of images containing that object or attribute. Hovering over an attribute node displays a bar chart of evaluation results.
@@
Figure 2. The ViperaBase prototype used in the formative study, showing the generated images (right) and a scene graph (left) for the user prompt (top). The scene graph is a node-link diagram where nodes indicate objects (or their attributes) within the images and edges indicate the semantic relationships. Bar charts will be shown when hovering on attribute nodes. :
Protocol. We first conducted semi-structured interviews to explore participants’ current auditing practices. Participants were then tasked with auditing the Stable Diffusion model using both prototypes in order (Baseline first, followed by ViperaBase). Each auditing session lasted five minutes, during which participants selected a prompt from the following prompts: a) A cinematic photo of a doctor, b) A family having a picnic in the park, c) A peaceful nature scene with diverse wildlife, and d) An award-winning chef preparing a gourmet meal. We designed these prompts to cover various topics (e.g., humanities, nature) and potential issues (e.g., bias on gender, profession, and race) while preventing users from the cold-start issue. The prompt would be used to generate 30 images, and the participants were asked to explore these images for as many insights as possible and note their insights and evidence in brief phrases or sentences. After interacting with each prototype, participants shared their reasoning processes. Finally, they provided feedback on their perceptions of both prototypes and requirements for Vipera. All study sessions were conducted online, and the average duration of each session was approximately 50 minutes.

### v2

@@ Abstract.
Despite their increasing capabilities, text-to-image generative AI systems are known to produce biased, offensive, and otherwise problematic outputs. While recent advancements have supported testing and auditing of generative AI, existing auditing methods still face challenges in supporting effective exploration of the vast space of AI-generated outputs in a structured way. To address this gap, we conducted formative studies with five AI auditors and synthesized five design goals for supporting systematic AI audits. Based on these insights, we developed Vipera, an interactive auditing interface that employs multiple visual cues, including a scene graph, to facilitate image sensemaking and inspire auditors to explore and hierarchically organize the auditing criteria. Additionally, Vipera leverages LLM-powered suggestions to enable exploration of unexplored auditing directions. Through a controlled experiment with 24 participants experienced in AI auditing, we demonstrate Vipera’s effectiveness in helping auditors navigate large AI output spaces and organize their analyses while engaging with diverse criteria.
AI Auditing, Responsible AI, Generative AI, Visual Analytics
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spainccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interactive systems and tools
@@ 1. Introduction
Generative text-to-image (T2I) models are gaining popularity for their ability to enhance the efficiency and expressiveness of creative activities. However, the associated risks are significant; these models can produce images that may perpetuate biases, cause offense, or disseminate misleading information (Naik and Nushi, 2023; Qi et al., 2024). To surface and address these issues, researchers and practitioners have turned to AI auditing—a process of repeatedly testing an algorithm with various inputs and observing the corresponding outputs—to better understand its behavior and potential external impacts. (Sandvig et al., 2014; Metaxa et al., 2021; Birhane et al., 2024).
@@
Recent research in HCI and responsible AI (RAI) has explored tools and processes to support diverse auditors in AI audits (Lam et al., 2022; Shen et al., 2021; Cabrera et al., 2021a; Deng et al., 2023a; Maldaner et al., 2024; Claire et al., 2024). However, auditing generative T2I models in a systematic and scaled way remains challenging. To start with, manual evaluation of these systems can be time-consuming, often limiting assessments to a small number of prompts or outputs, which can hardly uncover the comprehensive issues of the system. Developers often face challenges in exploring potentially productive auditing directions and gathering scaled data for each identified issue (Lam et al., 2022; Deng et al., 2025b). Reviewing and aggregating audit reports also remains a demanding task (Cabrera et al., 2021a; Deng et al., 2023a; Ojewale et al., 2025)
@@
To tackle these challenges, automated approaches have also been proposed, often involving AI-supported output labeling and applied to textual models (Shankar et al., 2024; Rastogi et al., 2023b; Arawjo et al., 2024). However, two primary obstacles persist when it comes to auditing T2I models. First, images encompass a broad range of semantics, resulting in a vast auditing space that is difficult to characterize with a concise list of criteria as often used for evaluating texts (Suh et al., 2024). For example, an image may feature multiple characters, each of whom can be evaluated based on their clothing, and the clothing can be further assessed in terms of its cultural nature and relationship to the character’s profession. To systematically explore this space, it is crucial to keep auditors aware of the unknown unknowns - criteria that are yet to be explored (Kiela et al., 2021; Deng et al., 2025b). Additionally, our observations from a formative study with five auditors suggest that many auditors often rely on intuition or personal experience during the auditing process, rather than adopting a formal approach. There is a lack of structured methodologies for navigating and exploring this extensive auditing space.
@@
In this paper, we propose Vipera [^1] Vipera stands for ”Visual Intelligence-Promoted End User Auditing”, an interactive system for streamlining and enhancing the systematicness of large-scale T2I model auditing. Vipera facilitates structured, multi-faceted analysis through the visual cue of a scene graph (Johnson et al., 2015), where auditing criteria are organized hierarchically, coordinated with visual statistics, and associated with the scenic semantics of images. Additionally, Vipera incorporates LLM-powered auditing suggestions to uncover new avenues for analysis and facilitate exploration within the auditing space. In particular, it provides audit analysis support by highlighting differences between images to suggest new criteria, and prompt suggestions to inspire exploration of new topics, with optional user-customizable keywords for targeted guidance. To support auditors in documenting and reporting their findings, Vipera allows users to bookmark visual evidence immediately upon getting insights, incorporates a dedicated note view for structured note construction, and supports LLM-powered note completion for better efficiency.
@@
Through a controlled user study with 20 general auditors and 4 expert auditors, we demonstrate Vipera’s effectiveness in helping auditors navigate the large auditing space and organize their analyses while engaging with diverse criteria. In particular, we discovered that the blended guidance prompted participants to explore more prompts, images, and auditing criteria compared to conditions without guidance (see Section 6.1.2). We also demonstrate that visual and LLM-driven guidance are highly complementary, and blending them can improve performance, reduce user workload on average, and encourage a more systematic auditing process (see Section 6.2.3). Furthermore, we have revealed several interesting patterns in their auditing process that serve as the basis for future personalized auditing tools (see Section 6.2.4). For example, we observed breadth-oriented and depth-oriented patterns in both the creation of prompts and criteria, suggesting a need for personalized auditing guidance.
@@
  • An empirical study on 24 participants with various backgrounds in auditing that evaluates the system’s usability as well as the distinct and combined effects of these guidance modalities, demonstrating that their integration leads to more systematic and thorough auditing with reduced cognitive load compared to three ablated versions of the system.
@@ 2.1. Auditing generative AI (at scale)
Generative AI (GenAI) systems have sparked increasing societal concerns due to potential harmful behaviors, such as social biases and violence (Diberardino et al., 2024). This has led to increased focus on detecting and mitigating these issues through benchmarks such as SafetyBench (Zhang et al., 2024) and AgentHarm (Andriushchenko et al., 2025). However, the limited diversity of inputs and contexts in these benchmarks hampers real-world performance assessments (Kiela et al., 2021).
@@
AI audits have gained prominence as a method for uncovering biased, discriminatory, or otherwise harmful behaviors in algorithmic systems (Noble, 2018; Asplund et al., 2020; Sweeney, 2013; Prates et al., 2020; Buolamwini and Gebru, 2018; Hannak et al., 2014; Sandvig et al., 2014; Metaxa et al., 2021; Birhane et al., 2024; Shen et al., 2022). More recent research has demonstrated the value of conducting AI audits to ensure safer and responsible GenAI, often even involving general end users in the auditing process, as they can uncover overlooked cases and provide insights (Shen et al., 2021; Lam et al., 2022; Cabrera et al., 2021b; Deng et al., 2025b). For example, Mack et al. engaged 25 people with disabilities to review and reflect on images generated by T2I systems. They uncovered several nuanced societal stereotypes that extended beyond existing taxonomies, such as “perpetuating broader narratives in society around disabled people as primarily using wheelchairs, being sad and lonely, incapable, and inactive” (Mack et al., 2024). Similarly, Shelby et al. conducted workshops with 15 cross-cultural artists to synthesize folk theories related to the use of T2I models, their potential harms, and harm reduction strategies perceived by artists (Shelby et al., 2024). Deng et al. created WeAudit system to engage 45 users in individually and collectively auditing GenAI, with scaffolds to support users in providing actionable insights to AI developers (Deng et al., 2025b). More recent work in HCI and RAI has also explored supporting auditing with different audiences, such as youth (Solyst et al., 2025), as well as in different contexts, such as medical domains (Morrison et al., 2025). These tools primarily focus on engaging people without prior auditing experience and enabling them to contribute mostly qualitative analyses of generative AI systems.
@@
Prior research has primarily engaged people in auditing limited sets of GenAI outputs, while acknowledging the need for better toolings for scalable, systematic auditing (Shelby et al., 2024; Deng et al., 2025b). However, research in HCI shows that crowdsourced auditing remains challenging and costly due to data scale and often narrow evaluation criteria (Cabrera et al., 2021a; Deng et al., 2023a). To this end, recent studies have begun to focus on scalable auditing of generative AI, primarily addressing two key challenges. First, evaluating large numbers of outputs within limited timeframes: here, human-in-the-loop auditing combines large language models for automatic evaluation with user-facing summaries (Zheng et al., 2024; Shankar et al., 2024; Rastogi et al., 2023a), and annotations and visualizations have been introduced to support interpretation (Gero et al., 2024; Deng et al., 2025a). Second, early auditing tools often relied on rigid quantitative metrics, but recent efforts allow users to define their own criteria (Kim et al., 2024).
@@
However, the aforementioned work on improving systematic auditing has mainly focused on auditing text-to-text models. Extending them to text-to-image models introduces new challenges: Images contain rich semantic information (e.g., objects and styles) that can lead to a diverse range of auditing criteria, yet there is a notable lack of tools to assist auditors in interpreting auditing results for iterative assessments in the image context (Pettersson, 1993). Despite recent research effort from the AI safety and CV community on automated red-teaming to reveal risks (Li et al., 2024a; D’Incà et al., 2024) and designing harm-resistant generative models to mitigate them (Li et al., 2024b; Schramowski et al., 2023; Orgad et al., 2023), human-driven and human-AI collaborative auditing is still deemed crucial for revealing subtle, contextual harmful behavior patterns closely related to domain knowledge and live experiences, urging for mixed-initative, systematic auditing tools (Rastogi et al., 2023b; Cabrera et al., 2023; Lam et al., 2022). To fill this gap, our work introduces Vipera, which integrates visual cues and AI-driven auditing suggestions, to support structured and systematic exploration of auditing criteria in the image domain. In contrast to prior auditing tools, Vipera enables both breadth (rapid exploration across many criteria) and depth (guided reasoning with AI-generated insights), helping auditors move beyond surface-level observations toward more comprehensive and reproducible image audits by organically blending visual and AI-driven supports. In the next section, we expand on the unique challenges of auditing image context.
@@ 2.2. Visual analytics for sensemaking image collection
Sensemaking of large image collections plays a key role in auditing T2I models at scale, and visual analytics approaches have proven effective in this context (Afzal et al., 2023). One common method involves clustering images based on pixel or vector embeddings for visualization, which supports various analytical tasks such as search and exploration, comparison, and visual summarization (Zahálka et al., 2020; Schmidt et al., 2013; Pan et al., 2019). To incorporate image semantics alongside visual features, recent works identify semantic objects within images and extract concise textual representations—such as keywords, descriptions, or captions. These text analysis results facilitate tasks like semantic categorization and pattern mining (Xie et al., 2018; Yee et al., 2003; Li et al., 2024c). However, characterizing the relationships between objects using plain text can be a challenging task. To address this, scene graphs have emerged as a more expressive representation, where nodes denote semantic objects and edges illustrate their relationships. This representation enhances detailed understanding of image collections and improves analysis performance (Fang et al., 2017; Johnson et al., 2015).
With the advent of generative T2I models, a recently emerged task is to visualize the relationship between prompts and corresponding images. For instance, DreamSheets organizes prompts and images in a spreadsheet layout to facilitate comparison (Almeda et al., 2024). Similarly, PrompTHis introduces an image variant graph to illustrate the semantic transitions of image clusters as users create new prompts (Guo et al., 2024). However, these approaches often struggle to support sensemaking in large-scale image sets, where variations among outputs for a single prompt can lead to overly condensed visual clusters.
Drawing from prior sensemaking work and situating it in the context of AI auditing, Vipera extends prior AI auditing interfaces by integrating several key forms of visual support to enable more effective and systematic auditing practices. In particular, Vipera first embeds bar charts within a tree-based scene graph, revealing distributional patterns for images related to individual prompts and allowing for comparisons across prompts. Besides, Vipera leverages the scene graph as a visual aid to foster systematic and structured AI auditing while providing inspiration to auditors. Additionally, Vipera incorporates various view coordinations to facilitate navigation between auditing results and raw images for enhanced sensemaking. Our work further contributes by evaluating these visualizations in the context of systematic AI auditing, offering critical insights for developing more effective auditing tools for text-to-image systems.
@@ 3.1.1. Study setup
Prototypes. Following parallel prototyping (Dow et al., 2010), we designed two prototypes for the formative study, referred to as ViperaBase and Baseline in the following discussions. As illustrated in Figure 1, ViperaBase features a scene graph view, i.e., the zoomable node-link diagram, alongside a prompt input box and image view. In this view, blue nodes represent objects, and green nodes indicate their attributes, with labeled edges illustrating relationships. The size of each node reflects the number of images containing that object or attribute. Hovering over an attribute node displays a bar chart of evaluation results. By contrast, the Baseline prototype includes all views of ViperaBase except the scene graph view, i.e., user can audit the model only by inspecting the images and trying new prompts. All images were generated by Stable Diffusion v3-medium, while the scene graph was created by aggregating the scene graphs of each image generated by LLaVA v1.5.
@@
Figure 1. The ViperaBase prototype used in the formative study, showing the generated images (right) and a scene graph (left) for the user prompt (top). The scene graph is a node-link diagram where nodes represent objects (or their attributes) within the images and edges represent the semantic relationships. Bar charts will be shown when hovering over attribute nodes.
The image displays the ViperaBase prototype where the user is auditing a T2I model with the prompt ”a cinematic photo of a doctor” (top). On the left, a scene graph in the network layout serves as a structural representation of the semantics within the multiple images on the right generated by the model. The nodes within it represent semantic objects (e.g., doctor; in green color) or the attributes of these objects (e.g., doctor_profession; in cyan color). The node size encodes the frequency of related terms, with ”doctor_profession” being the most prominent. On the right, a collage features various images of male and female doctors in clinical settings, each wearing medical attire and stethoscopes.
Protocol. We first conducted semi-structured interviews to explore participants’ current auditing practices. Participants were then tasked with auditing the Stable Diffusion model using both prototypes in order (Baseline first, followed by ViperaBase). Each auditing session lasted five minutes, during which participants selected a prompt from the following prompts: a) A cinematic photo of a doctor, b) A family having a picnic in the park, c) A peaceful nature scene with diverse wildlife, and d) An award-winning chef preparing a gourmet meal. We designed these prompts to cover various topics (e.g., humanities, nature) and potential issues (e.g., bias related to gender, profession, and race) while preventing users from the cold-start issue. The prompt would be used to generate 30 images, and participants were asked to explore these images for as many insights as possible, noting their insights and evidence in brief phrases or sentences. After interacting with each prototype, participants shared their reasoning processes. Finally, they provided feedback on their perceptions of both prototypes and requirements for Vipera. All study sessions were conducted online, with an average duration of approximately 50 minutes for each session.